A rigid plank 5.00 m long has mass 30.0 kg and supports a 75.0 kg block 2.00 m from the left end. How much force is needed to support each end of the plank?

1
Expert's answer
2022-03-06T18:06:34-0500
"mg*2+Mg*2.5=R_B*5""R_B=441\:\rm N""R_A=mg+Mg-R_B=588\:\rm N"